# Long-tailed Adversarial Training with Self-Distillation

### v1

@@ 3.2 Theoretical Analysis
Let the long-tailed dataset for a binary classification task data $\mathcal{S}$ with imbalance ratio $r\geq 1$, i.e., the ratio of the number of instances in the head class ($y=+1$) to the number of instances in the tail class ($y=-1$) is $r$. We assume Gaussian mixture distribution, which is similar to Xu et al. (2021); Lee et al. (2024) as follows.
@@
For a data distribution $\mathcal{S}$, the optimal natural classifier $f^{*}_{nat}$ and robust classifier $f^{*}_{rob}$ exhibit the following standard and robust errors for the tail class $-1$ with perturbation margin $0<\epsilon<\eta$, respectively:
@@ 3.3 Empirical Analysis
Figure 3 shows the test accuracy of the tail class ($y=-1$) of the logistic regressions. The results indicate a clear trend: as the imbalance ratio grows, the test accuracy for the tail class drops across both natural and adversarial training scenarios. More notably, adversarial training consistently produces more robust errors than natural training. These results align with our theoretical predictions: imbalanced data amplifies clean and robust errors for the tail class, and adversarial training further exacerbates robust errors.
@@
Input: Long-tailed dataset $\mathcal{D}$, batch size $N$, epochs $T$, learning rate $\mu$, hyperparameters $\alpha$, $\gamma$,
@@
                               $+\alpha\cdot\mathcal{L}_{KD}(f(\bm{x}^{\prime}_{i}),f_{B}(\bm{x}_{i}))\Big{]}$            # Self-distillation
@@ 5.3.3 Effect of Balanced Subset
[…]

### v3

@@ 3.2 Theoretical Analysis
Let the long-tailed dataset for a binary classification task data $\mathcal{S}$ with imbalance ratio $r\geq 1$, i.e., the ratio of the number of instances in the head class ($y=+1$) to the number of instances in the tail class (y=1y=-1) is $r$. We assume Gaussian mixture distribution, which is similar to Xu et al. (2021); Lee et al. (2024) as follows.
@@
For a data distribution $\mathcal{S}$, the optimal natural classifier $f^{*}_{nat}$ and robust classifier $f^{*}_{rob}$ exhibit the following standard and robust errors for the tail class 1-1 with perturbation margin $0<\epsilon<\eta$, respectively:
@@ 3.3 Empirical Analysis
Figure 3 shows the test accuracy of the tail class (y=1y=-1) of the logistic regressions. The results indicate a clear trend: as the imbalance ratio grows, the test accuracy for the tail class drops across both natural and adversarial training scenarios. More notably, adversarial training consistently produces more robust errors than natural training. These results align with our theoretical predictions: imbalanced data amplifies clean and robust errors for the tail class, and adversarial training further exacerbates robust errors.
@@
Input: Long-tailed dataset $\mathcal{D}$, batch size NN, epochs $T$, learning rate $\mu$, hyperparameters $\alpha$, $\gamma$,
@@
             $\alpha\cdot\mathcal{L}_{KD}(f(\bm{x}^{\prime}_{i}),f_{B}(\bm{x}_{i}))\Big{]}$      # Self-distillation
@@ 5.3.3 Effect of Balanced Subset
To evaluate the effect of the balanced subset $D_{B}$, we compare the performance of models trained with PGD-AT on $D_{B}$ against Robal, Reat, and BSL trained on $D$, which incorporate techniques like balanced softmax to address long-tailed distributions. For simplicity, we denote $\text{PGD-AT}_{D_{B}}$ as the model trained with PGD-AT on the balanced subset $D_{B}$.
Figure 5: Clean and robust accuracy according to different teachers. Robal, Reat, and Bsl were trained with 100 epochs, while $\text{PGD-AT}_{D_{B}}$ used a teacher trained with 30 epochs.
As shown in Table 6, while the PGD-AT model trained on $D_{B}$ achieved the lowest overall performance, it demonstrated the best results for tail classes, T-Clean and T-PGD. This suggests that even with fewer training epochs, the balanced subset is effective for improving performance on tail classes. In Figure 5, we apply the same methods with different teachers where the performance is summarized in Table 6. Interestingly, the best results were achieved when we utilize $\text{PGD-AT}_{D_{B}}$ as the teacher model, despite having the lowest overall performance. Specifically, the performance on the tail classes highlights the effectiveness of the teacher trained on the balanced subset, as it demonstrates superior performance on the tail class compared to other methods. This underscores the utility of the balanced subset in improving tail class performance. Additionally, although we trained the teacher for self-distillation using a simple method, PGD-AT, developing a more effective teacher remains an area for future work.